Solve the polynomial equation by factoring and then using the zero product principle

2x to the 4th power = 128x

Answer by srinivas(5)

(2x)^4=128*x
16x^4=128*x
16x^4-128x=0
x(16x^3-128)=0
x=0 or 16x^3-128=0
16x^3=128
x^3=8
x=2
therefore x=0,2